# Characteristic Learning for Provable One Step Generation

### v3

@@ Abstract.
We propose the characteristic generator, a novel one-step generative model that combines the efficiency of sampling in Generative Adversarial Networks (GANs) with the stable performance of flow-based models. Our model is driven by characteristics, along which the probability density transport can be described by ordinary differential equations (ODEs). Specifically, We estimate the velocity field through nonparametric regression and utilize Euler method to solve the probability flow ODE, generating a series of discrete approximations to the characteristics. We then use a deep neural network to fit these characteristics, ensuring a one-step mapping that effectively pushes the prior distribution towards the target distribution. In the theoretical aspect, we analyze the errors in velocity matching, Euler discretization, and characteristic fitting to establish a non-asymptotic convergence rate for the characteristic generator in $2$-Wasserstein distance. To the best of our knowledge, this is the first thorough analysis for simulation-free one step generative models. Additionally, our analysis refines the error analysis of flow-based generative models in prior works. We apply our method on both synthetic and real datasets, and the results demonstrate that the characteristic generator achieves high generation quality with just a single evaluation of neural network.
@@ 1. Introduction
Is it possible to develop a one-step generative model that combines the efficient sampling of GANs with the stable performance of diffusion or flow-based models? If so, can we establish a rigorous error analysis for this generative model?
Several recent papers have made progress in addressing the first question using various techniques such as distillation (Luhman and Luhman, 2021; Salimans and Ho, 2022; Song et al., 2023; Zhou et al., 2024), operator learning (Zheng et al., 2023a), or trajectory models (Kim et al., 2024; Ren et al., 2024). For a more detailed discussion, please refer to Section 5.1. Despite these recent advancements, a unifying mathematical framework for designing and analyzing the one-step generative models remains largely limited (Li et al., 2024b). This paper aims to fill this gap and provide a positive answer to the aforementioned questions. Specifically, we introduce a comprehensive framework, known as the characteristic generator, aiming to streamline the sampling process using a single evaluation of the neural network. Our model merges the sampling efficiency of GANs with the promising performance of flow-based models. Furthermore, we present a rigorous error analysis for the characteristic generator. Through numerical experiments, we validate that our approach generates high-quality samples that is comparable to those generated through flow-based models, all while requiring just a single evaluation of characteristic generator.
@@ 1.1. Contributions
    We introduce a one-step generative model called the characteristic generator, which efficiently pushes a Gaussian distribution to the target distribution without the need for simulation. We present a probability density transport equation via stochastic interpolants, resulting in probability flow ODEs along characteristics. The velocity field of the ODE is estimated through a least-squares regression. Subsequently, we generate a sequence of discrete approximations to characteristics by numerically solving the ODE. Finally, we train the characteristic generator by fitting these characteristics using a deep neural network.
@@
    We provide a rigorous error analysis for the characteristic generator. Specifically, we derive a convergence rate ${\mathcal{O}}(n^{-\frac{1}{d+3}})$ for velocity matching (Theorem 3.9), which improves the rates in previous works such as Chen et al. (2023b). Additionally, we propose an error bound for the $2$-Wasserstein distance between the data generated by the Euler method and the target distribution (Theorem 3.11), which is of independent interest. Lastly, we present a non-asymptotic convergence rate for the characteristic generator in the $2$-Wasserstein distance, demonstrating that it achieves a rate same to that of Euler sampling (Theorem 3.13). To the best of our knowledge, this is the first analysis conducted for one-step sampling of flow-based models, providing valuable theoretical insights for distillation (Salimans and Ho, 2022; Song et al., 2023), operator learning (Zheng et al., 2023a), or trajectory model (Kim et al., 2024; Ren et al., 2024).
@@
    We validate the generation quality and sampling efficiency of the characteristic generator using both synthetic and real data through numerical experiments. The experiment results demonstrate that our method can generate samples of high quality from noise with just one neural network evaluation. In comparison to previous one-step generative models without the aid of GANs (Salimans and Ho, 2022; Song et al., 2023; Kim et al., 2024) , our characteristic generator significantly improves the generation quality in CIFAR-10, as shown in Table 4. Furthermore, with only a few iterations, our characteristic generator achieves comparable generation performance to the state-of-the-art method (Kim et al., 2024, CTM), without the need for additional GAN training as employed by CTM.
@@ 1.2. Main Results
Let $\mu_{1}\in P_{\operatorname{ac}}({\mathbb{R}}^{d})$ be a target probability distribution with $\,\mathrm{d}\mu_{1}(x)=\rho_{1}(x)\,\mathrm{d}x$. In the context of generative learning, one only has access to data samples from $\mu_{1}$ but have no access to the density function $\rho_{1}$ itself. Let $\mu_{0}\in P_{\operatorname{ac}}({\mathbb{R}}^{d})$ be a known prior distribution. Suppose that there exists a smooth mapping $G^{*}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$, which pushes-forward the prior distribution $\mu_{0}$ onto the target distribution $\mu_{1}$, that is,
@@
The equation (1.1) is known as the normalizing equation (Rozen et al., 2021). The goal of the generative learning, at least formally, is to find an estimator $\widehat{G}$ of the push-forward operator $G^{*}$ based on finite samples drawn from $\mu_{1}$.
In this work, we aim to construct the desired push-forward operator via a probability flow ODE
$\,\mathrm{d}x(t)=b^{*}(t,x(t))\,\mathrm{d}t,\quad x(0)=x_{0}\sim\mu_{0},$
where the velocity field $b^{*}$ is given as (2.3). Denote by $\mu_{t}$ the distribution of $x(t)$ for each $t\in(0,1)$, and define the probability flow $g_{t,s}^{*}$ as $g_{t,s}^{*}(x(t))=x(s)$ for each $10\leq t\leq s<$ 1. Let $\widehat{b}$ be the estimated velocity field obtained by velocity matching (2.8), and let $\widehat{E}_{0,K}^{\tau}$ be the numerical approximation to the ODE solution by Euler method (2.10). Denote by $\widehat{g}_{s,t}$ be the estimation of the flow $g_{t,s}^{*}$ defined as (2.12), which is referred to the characteristic generator.
@@
Assumption 2.
The target distribution $\mu_{1}=N(0,\sigma^{2}I_{d})*\nu$, where $\nu$ is a distribution with $\operatorname{supp}(\nu)\subseteq[0,1]^{d}$.
Assumption 2 requires the target distribution to be smoothed by a Gaussian distribution. This assumption is essential as it ensures desirable properties of the probability flow ODE, such as bounded moments and the Lipschitz property of the velocity field. It is noteworthy that this assumption can be considered relatively mild, given that the smoothed distribution $\mu_{1}$ is an approximation of the original distribution $\nu$, particularly when the variance $\sigma^{2}$ of the Gaussian distribution is small. Further details can be found in Sections 3.1 and 3.2.
@@
Suppose that Assumptions 1 and 2 hold. Let ${\mathcal{S}}$ be a set of $n$ samples independently and identically drawn from the target distribution $\mu_{1}$. Then it follows that
@@
Suppose that Assumptions 1 and 2 hold. Let ${\mathcal{S}}$ be a set of $n$ samples independently and identically drawn from the target distribution $\mu_{1}$. By setting a sufficiently large number of time steps $K$ for Euler method, the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}W_{2}^{2}\Big{(}(\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{1}\Big{)}\Big{]}\leq C_{T}^{1}n^{-\frac{2}{d+3}}\log^{2}n+C_{T}^{2}W_{2}^{2}\big{(}\mu_{0},\mu_{1}\big{)},$
where $C_{T}^{1}$ and $C_{T}^{2}$ are two constants depending on $d$, $\sigma$ and $T$. Further, as the stopping time $T\rightarrow 1$, the constant $C_{T}^{1}$ tends to infinity while $C_{T}^{2}$ decreases to zero.
@@
Suppose that Assumptions 1 and 2 hold. Let ${\mathcal{S}}$ be a set of $n$ samples independently and identically drawn from the target distribution $\mu_{1}$, and let ${\mathcal{Z}}$ be the set of numerical solutions by Euler method with a sufficiently large number. Then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}W_{2}^{2}\Big{(}(\widehat{g}_{t,s})_{\sharp}\mu_{t},\mu_{s}\Big{)}\,\mathrm{d}s\mathrm{d}t\Big{]}\leq C_{T}n^{-\frac{2}{d+3}}\log^{2}n,$
where $C$ is a constant depending on $d$, $\sigma$ and $T$.
@@
Probability flow ODE Convergence rate
Linear interpolants ${\mathcal{O}}(n^{-\frac{2}{3(d+3)}})$ Corollary 3.14
Föllmer flow ${\mathcal{O}}(n^{-\frac{2}{5(d+3)}})$ Corollary 3.15
@@ 1.3.1. Wasserstein Distance
(1.2) $W_{2}(\mu_{0},\mu_{1})=\inf\Big{\{}{\mathbb{E}}^{1/2}\big{[}\|X_{0}-X_{1}\|_{2}^{2}\big{]}:\leavevmode\nobreak\ \operatorname{law}(X_{0})=\mu_{0},\leavevmode\nobreak\ \operatorname{law}(X_{1})=\mu_{1}\Big{\}}.$
@@ 1.3.3. Notations
[…]

### v5

@@ Abstract.
We propose the characteristic generator, a novel one-step generative model that combines the efficiency of sampling in Generative Adversarial Networks (GANs) with the stable performance of flow-based models. Our model is driven by characteristics, along which the probability density transport can be described by ordinary differential equations (ODEs). Specifically, we first estimate the underlying velocity field and use the Euler method to solve the probability flow ODE, generating discrete approximations of the characteristics. A deep neural network is then trained to fit these characteristics, creating a one-step map that pushes a simple Gaussian distribution to the target distribution. In the theoretical aspect, we provide a comprehensive analysis of the errors arising from velocity matching, Euler discretization, and characteristic fitting to establish a non-asymptotic convergence rate in the 2-Wasserstein distance under mild data assumptions. Crucially, we demonstrate that under a standard manifold assumption, this convergence rate depends only on the intrinsic dimension of data rather than the much larger ambient dimension, proving our model’s ability to mitigate the curse of dimensionality. To our knowledge, this is the first rigorous convergence analysis for a flow-based one-step generative model. Experiments on both synthetic and real-world datasets demonstrate that the characteristic generator achieves high-quality and high-resolution sample generation with the efficiency of just a single neural network evaluation.
@@ 1. Introduction
How can we develop a one-step generative model that combines the efficient sampling of GANs with the stable performance of diffusion or flow-based models? How can we establish a rigorous error analysis for this generative model?
Several recent papers have made progress in addressing the first question using various techniques such as distillation (Luhman and Luhman, 2021; Salimans and Ho, 2022; Song et al., 2023; Zhou et al., 2024), operator learning (Zheng et al., 2023a), or trajectory models (Kim et al., 2024; Ren et al., 2024). For a more detailed discussion, please refer to Section 5.1. Despite these recent advancements, a unifying mathematical framework for designing and analyzing the one-step generative models remains largely limited (Li et al., 2024b). This paper aims to fill this gap and provide an answer to the aforementioned questions. Specifically, we introduce a comprehensive framework, known as the characteristic generator, aiming to streamline the sampling process using a single evaluation of the neural network. Our model merges the sampling efficiency of GANs with the promising performance of flow-based models. Furthermore, we present a rigorous error analysis for the characteristic generator. Through numerical experiments, we validate that our approach generates high-quality samples that is comparable to those generated through flow-based models, all while requiring just a single evaluation of the neural network.
@@ 1.1. Contributions
    We introduce the “characteristic generator,” a one-step generative model that forms the basis of a unified mathematical framework for recent flow-based distillation methods. Our approach involves modeling the probability density transport equation with stochastic interpolants, which yields a probability flow ODE. By learning to approximate the characteristic curves of this ODE, the generator can directly transform a prior distribution to the target distribution without iterative simulation.
@@
    We provide a rigorous error analysis for the characteristic generator. Specifically, we derive a convergence rate ${\mathcal{O}}(n^{-\frac{1}{d+3}})$ for velocity matching (Theorem 3.10). Additionally, we propose an error bound for the $2$-Wasserstein distance between the distribution of data generated by the Euler method and the target distribution (Theorem 3.12), which is of independent interest. Lastly, we present a non-asymptotic convergence rate for the characteristic generator in the $2$-Wasserstein distance (Theorem 3.14). These findings also provide valuable theoretical understanding for distillation (Salimans and Ho, 2022; Song et al., 2023), operator learning (Zheng et al., 2023a), or trajectory model (Kim et al., 2024; Ren et al., 2024).
@@
    We prove that the characteristic generator mitigates the curse of dimensionality (Corollaries 3.19 and 3.20) under a standard manifold assumption. We show the convergence rate depends on the data’s low intrinsic dimension rather than the high ambient dimension. To our knowledge, this is the first end-to-end analysis that formally explains the strong performance of flow-based one-step generative models on high-dimensional data.
  4. (iv)
    We conduct extensive experiments on synthetic and real-world data (CIFAR-10, CelebA-HQ), demonstrating that the characteristic generator produces high-quality samples in a single network evaluation. Our model significantly outperforms prior non-adversarial one-step models (Salimans and Ho, 2022; Song et al., 2023; Kim et al., 2024) on CIFAR-10. Furthermore, it achieves performance competitive with state-of-the-art methods (Kim et al., 2024, CTM) in just a few iterations, without requiring additional adversarial training. The successful application to high-resolution image generation ($256\times 256$ and $512\times 512$) further validates the scalability of our approach.
@@ 1.2. Main Results
Let $\mu_{0}\in P_{\operatorname{ac}}({\mathbb{R}}^{d})$ be a known prior distribution, and let $\mu_{1}\in P_{\operatorname{ac}}({\mathbb{R}}^{d})$ be the target distribution with an unknown density function $\rho_{1}(x)$. Suppose one has access to finite data samples from $\mu_{1}$. In generative learning, we aim to learn a push-forward map $G^{*}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ that transports the prior distribution $\mu_{0}$ onto the target distribution $\mu_{1}$. This relationship is defined by the normalizing equation (Rozen et al., 2021):
@@
Formally, the goal of the generative learning is to find an estimator $\widehat{G}$ of the push-forward operator $G^{*}$ based on finite samples drawn from $\mu_{1}$.
Remark 1.1 (Regularity of the push-forward map).
The learnability of the push-forward map $G^{*}$ requires its regularity. Without crucial properties like boundedness and Lipschitz continuity, the map can be ill-behaved, making it practically impossible for function approximators like neural networks to learn. In this work, we define $G^{*}$ via a probability flow ODE (1.2), and establish the necessary regularity conditions in Section 3.2.
In this work, we construct the desired push-forward operator via a probability flow ODE:
(1.2) $\mathrm{d}x(t)=b^{*}(t,x(t))\,\mathrm{d}t,\quad x(0)=x_{0}\sim\mu_{0},$
where the velocity field $b^{*}$ is given as (2.3), and $x(t)$ is knwon as the characteristic curve of the transport equation (2.2). Denote by $\mu_{t}$ the distribution of $x(t)$ for each $t\in(0,1)$, and define the flow $g_{t,s}^{*}$ as $g_{t,s}^{*}(x(t))=x(s)$ for each $0\leq t\leq s\leq 1$, which transports particles along the characteristic curves. It is apparent that $(g_{t,s}^{*})_{\sharp}\mu_{t}=\mu_{s}$. Then our target push-forward map is defined as $G^{*}:=g_{0,1}^{*}$. Let $\widehat{b}$ be the estimated velocity field obtained by velocity matching (2.8), and let $\widehat{E}_{0,K}^{\tau}$ be the numerical approximation to the ODE solution by Euler method (2.9). Denote by $\widehat{g}_{s,t}$ be an estimation of $g_{t,s}^{*}$ defined as (2.11), which is referred to the characteristic generator.
@@
Assumption 2 (Target distribution).
There exists an unknown constant $\sigma>0$, such that
$\mu_{1}(x)=N(0,\sigma^{2}I_{d})*\nu:=\int\varphi_{d}\Big(\frac{x-x^{\prime}}{\sigma}\Big)\,\mathrm{d}\nu(x^{\prime}),$
where $\varphi_{d}$ represents the density of the $d$-dimensional standard Gaussian distribution, and $\mathrm{d}\nu(x)=p(x)\,\mathrm{d}x$ with $\operatorname{supp}(\nu)\subseteq[0,1]^{d}$.
Assumption 2 requires the target distribution to be a Gaussian convolution. In other words, for each random variable $X\sim\mu_{1}$, there exist two independent random variables $Z\in[0,1]^{d}$ and $W\sim N(0,I_{d})$, such that $X\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}Z+\sigma W$. This assumption is essential as it ensures desirable properties of the probability flow ODE, such as bounded moments and the Lipschitz property of the velocity field. Further details can be found in Sections 3.1 and 3.2. It is noteworthy that this assumption can be considered relatively mild, given that the smoothed distribution $\mu_{1}$ is an approximation of the original distribution $\nu$, particularly when the variance $\sigma^{2}$ of the Gaussian distribution is small. In addition, it encompasses non-log-concave distributions, such as Gaussian mixtures (Grenioux et al., 2024, Appendix C). Similar assumptions have been considered by Saremi et al. (2024); Grenioux et al. (2024); Beyler and Bach (2025); however, Beyler and Bach (2025) requires prior knowledge of the variance $\sigma^{2}$.
@@
Suppose that Assumptions 1 and 2 hold. Let $\mathcal{S}$} be a set of $n$ samples independently and identically drawn from the target distribution $\mu_{1}$. Suppose the depth and width of the velocity neural network are set properly. Then it follows that
@@
Suppose that Assumptions 1 and 2 hold. Let $\mathcal{S}$} be a set of $n$ samples independently and identically drawn from the target distribution $\mu_{1}$. Suppose the depth and width of the velocity neural network are set properly. Let $K$ be the number of time steps of Euler method. Then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}\Big[W_{2}^{2}\Big((\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{1}\Big)\Big]\leq C_{T}^{1}\Big\{n^{-\frac{2}{d+3}}\log^{2}n+\frac{\log n}{K^{2}}\Big\}+C_{T}^{2}W_{2}^{2}\big(\mu_{0},\mu_{1}\big),$
where $C_{T}^{1}$ and $C_{T}^{2}$ are two constants depending on $d$, $\sigma$ and $T$. Further, as the stopping time $T\rightarrow 1$, the constant $C_{T}^{1}$ tends to infinity polynomially while $C_{T}^{2}$ decreases to zero polynomially.
@@
Suppose that Assumptions 1 and 2 hold. Let $\mathcal{S}$} be a set of $n$ samples independently and identically drawn from the target distribution $\mu_{1}$, and let $\mathcal{Z}$} be the set of numerical solutions by Euler method with a sufficiently large number. Suppose the depth and width of the velocity neural network and the characteristic neural network are set properly, respectively. Let $m$ be the number of time steps of Euler method, and $m$ be the number of samples for characteristic fitting. Then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\Big[\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}W_{2}^{2}\Big((\widehat{g}_{t,s})_{\sharp}\mu_{t},\mu_{s}\Big)\,\mathrm{d}s\mathrm{d}t\Big]$
$\leq C_{T}^{1}\Big\{n^{-\frac{2}{d+3}}\log^{2}n+\frac{\log n}{K^{2}}\Big\}+C\Big\{m^{-\frac{2}{d+4}}\log^{2}m+\frac{\log m}{K}\Big\},$
where $C_{T}^{1}$ is a constant depending on $d$, $\sigma$ and $T$, and $C$ is a constant depending on $d$ and $\sigma$. Further, as the stopping time $T\rightarrow 1$, the constant $C_{T}^{1}$ tends to infinity.
@@
Probability flow ODE Assumption Convergence rate
Linear interpolants Assumptions 1 and 2 ${\mathcal{O}}(n^{-\frac{2}{3(d+3)}})$ Corollary 3.16
Föllmer flow Assumptions 1 and 2 ${\mathcal{O}}(n^{-\frac{2}{5(d+3)}})$ Corollary 3.17
The convergence rates in Table 1 suffer from the curse of dimensionality. However, in many practical applications, high-dimensional data – such as images or text documents – often lie close to a low-dimensional manifold embedded in the high-dimensional ambient space (Goodfellow et al., 2016; Bortoli, 2022; Jiao et al., 2023a; Oko et al., 2023). This observation motivates a refined assumption that captures the intrinsic low-dimensional structure of real-world data. We formalize this idea with the following assumption.
Assumption 3.
There exists an unknown constant $\sigma>0$ and $d^{*}\ll d$, such that
[…]
where $P\in{\mathbb{R}}^{d\times d^{*}}$ is a matrix whose column vectors are orthonormal in $\mathbb{R}^{d}$, and $\tilde{\nu}$ is a distribution with $\operatorname{supp}(\tilde{\nu})\subseteq[0,1]^{d^{*}}$.
We note that the distribution $\tilde{\nu}$ in Assumption 3 is supported on a subset of a $d^{*}$-dimensional hypercube. Consequently, its push-forward $\nu:=P_{\sharp}\tilde{\nu}$ lies on a $d^{*}$-dimensional manifold embedded in the ambient space $\mathbb{R}^{d}$. Although Gaussian convolution spreads $\nu$ throughout the full ambient space, the resulting distribution $\mu_{1}$ remains concentrated around the manifold, its density decaying exponentially with the distance from it. This manifold assumption is crucial, as it allows us to show that the convergence rate depends only on the intrinsic dimension $d^{*}$ rather than the ambient dimension $d$, thereby effectively mitigating the curse of dimensionality. The convergence rates of characteristic generators under manifold assumption are summarized in Table 2.
Table 2. Convergence rates of characteristic generator under manifold assumption.
Probability flow ODE Assumption Convergence rate
Linear interpolants Assumptions 1 and 3 ${\mathcal{O}}(n^{-\frac{2}{3(d^{*}+3)}})$ Corollary 3.19
Föllmer flow Assumptions 1 and 3 ${\mathcal{O}}(n^{-\frac{2}{5(d^{*}+3)}})$ Corollary 3.20
@@ 1.3.1. Wasserstein Distance
(1.3) $W_{2}(\mu_{0},\mu_{1})=\inf\Big\{{\mathbb{E}}^{1/2}\big[\|X_{0}-X_{1}\|_{2}^{2}\big]:\penalty 10000\ \operatorname{law}(X_{0})=\mu_{0},\penalty 10000\ \operatorname{law}(X_{1})=\mu_{1}\Big\}.$
@@ 1.3.3. Notations
The set of positive integers is denoted by ${\mathbb{N}}=\{1,2,\ldots\}$. We also denote ${\mathbb{N}}_{0}=\{0\}\cup{\mathbb{N}}_{+}$ for convenience. For a vector $x=(x_{1},\ldots,x_{d})\in{\mathbb{R}}^{d}$, we define its $\ell_{p}$-norms as $\|x\|_{p}^{p}=\sum_{i=1}^{d}|x_{i}|^{p}$ for $1\leq p<\infty$, with $\|x\|_{\infty}=\max_{1\leq i\leq d}|x_{i}|$. Denote by $\langle\cdot,\cdot\rangle$ the inner product between vectors, that is, $\langle x,y\rangle=\sum_{k=1}^{d}x_{k}y_{k}$, where $y=(y_{1},\ldots,y_{d})$. For a matrix $A\in{\mathbb{R}}^{m\times n}$, the operator norm induced by the $\ell_{2}$ vector norm is defined as $\|A\|_{\operatorname{op}}=\sup_{\|x\|_{2}=1}\|Ax\|_{2}$. Additionally, denote by ${\mathbb{B}}_{R}^{\infty}$ the $\ell_{\infty}$ ball in ${\mathbb{R}}^{d}$ with radius $R$, that is, ${\mathbb{B}}_{R}^{\infty}=\{x\in{\mathbb{R}}^{d}:\|x\|_{\infty}\leq R\}$. For a matrix $M$, we say $M\succeq 0$ if and only if it is positive definite. Let $A$ and $B$ be two matrices, denote $A\succeq B$ if and only if $(A-B)\succeq 0$. For a function $u(t)$ of time $t$, the time derivative is denoted by $\dot{u}$ or $\partial_{t}u$. Further, let $\ddot{u}$ denote the second-order time derivative. Additionally, we use $\nabla$ and $\nabla\cdot$ to denote the spatial gradient and divergence operators, respectively.